Statistics.

2- It has been reported that 9% of U.S. households do not own a car, 33% own one car, 38% own 2 cars, and 20% own 3 or more cShow your work.

Please be as neat as possible in answering the questions
Answer #1

This is a goodness of fit test for the given proportions against the observed frequencies. The expected frequencies for each of the value of number of cars owned are computed here as:

E(0) = 100*0.09 = 9
E(1) = 0.33*100 = 33
E(2) = 0.38*100 = 38
E(3 or more) = 0.2*100 = 20

The chi square test statistic here is computed as:

\chi^2 = \sum \frac{(O_i - E_i)^2}{E_i}

\chi^2 = \frac{11^2}{9} + \frac{2^2}{33} + \frac{15^2}{38} + \frac{2^2}{20}

\chi^2 = 19.6867

For k - 1 = 3 degrees of freedom, the p-value here is computed from the chi square distribution tables as:

p = P(\chi^2_3 > 19.6867) = 0.0002

As the given p-value here is 0.0002 which is very very low, therefore the test is significant here and we can reject the null hypothesis conclude that we have sufficient evidence that the car ownership in this community differs from the nation.